OpenStudy (jazzyfa30):

f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5 h[f(5)] - g[h(1)]

OpenStudy (primeralph):

Just plug in.

OpenStudy (primeralph):

@Jhannybean

OpenStudy (jazzyfa30):

ummmmm

OpenStudy (jhannybean):

first find f(5).\[\large f(5) = (5) +5 = \ ?\]

OpenStudy (jazzyfa30):

10

OpenStudy (jhannybean):

Then \[g(x) = x^2 -4x+5\]\[h(x) = 3x^2-2x+5\]Find h(1).\[h(1) = 3(1)^2 -2(1)+5 = \ ?\]

OpenStudy (jhannybean):

What do you get @jazzyfa30 ?

OpenStudy (jazzyfa30):

h=2????

OpenStudy (jhannybean):

3-2+5..= ? ...

OpenStudy (jazzyfa30):

oh 6

OpenStudy (jhannybean):

Yep. so.. if g(x) = x^2-4x+5, then g(h(1)) = g(6) = (6)^2 -4(6)+5 =...?

OpenStudy (jazzyfa30):

17???

OpenStudy (jhannybean):

I've got to head out in a few minutes :(

OpenStudy (jazzyfa30):

ok this is my last Q anyway

OpenStudy (jhannybean):

Yes, 17.

OpenStudy (jhannybean):

so we've got h(10-17)

OpenStudy (jhannybean):

that means we're solving for h(-7)

OpenStudy (jhannybean):

h(-7) = 3(-7)^2 - 2(-7) + 5 = ?

OpenStudy (jazzyfa30):

432????

OpenStudy (jhannybean):

147 +14 + 5 =?

OpenStudy (jazzyfa30):

ohh 166

OpenStudy (jhannybean):

yes good job.

OpenStudy (jazzyfa30):

so the answer is 166

OpenStudy (jhannybean):

Yes :)